OpenStudy (anonymous):

Al goes 8 mph; Ben goes 20 mph in the opposite direction. How many hours has Ben traveled when they are 180 miles apart?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

15 hours?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okayyy how did you get that answer?

OpenStudy (anonymous):

20x -8x = 180 12x = 180 x = 15 hrs

OpenStudy (anonymous):

except for it was wrong.. :/

OpenStudy (lilg132):

=O

OpenStudy (anonymous):

(45/7) hours

OpenStudy (lilg132):

8x + 20x = 180 28x = 180 28x / 28 = 180 / 28 x = 6.428571428571429 sorry :p

OpenStudy (anonymous):

6 hours and 42 minutes?

OpenStudy (lilg132):

6.42 hours

OpenStudy (anonymous):

it is actually 6 hours and 25.2 minutes..

OpenStudy (anonymous):

yeah i too took it in wrong sense....^^^^^ is right

OpenStudy (lilg132):

sorry my bad 0.42 * 60 = 25.2 minutes

OpenStudy (lilg132):

so the answer is 6 hours 25 minutes and 20 seconds